OpenStudy (anonymous):

Evaluate arccsc√2 a. pi/3 b. {pi/4±5pi/6} c. {pi/6, 5pi/6} d. {pi/3±2pi(n),5pi/3±2pi(n)}

OpenStudy (mertsj):

That is the same as saying that the sine of the angle is sqrt2/2

OpenStudy (anonymous):

Draw a \(\dfrac{\pi}{6},\dfrac{\pi}{3},\dfrac{\pi}{2}\) (30-60-90) triangle with labeled side lengths, and the answer will follow immediately by the definition of the cosecant function.

OpenStudy (mertsj):